OpenStudy (anonymous):

convert 86/125 to a percent? 68.8% is this correct

OpenStudy (callisto):

convert 86/125 to a percent = 86/125 x 100% = 68.8% = correct :)

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome!